OpenStudy (anonymous):

Variable c is 9 more than variable w. Variable c is also 3 less than variable w. Which pair of equations best models the relationship between c and w? c = w − 9 c = 4w + 3 c = w + 9 c = w − 3 w = c + 9 w = c − 3 w = c − 9 w = 4c + 3

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@fallenangelorchid

OpenStudy (mathstudent55):

Let's start with this. "Variable c is 9 more than variable w."

OpenStudy (mathstudent55):

That means c = w + 9 If we subtract 9 from both sides, we get c - 9 = w That means the first statement can be either c = w + 9, or c - 9 = w From here, we see that choices 2 and 4 have the correct first equation. We can eliminate choices 1 and 3.

OpenStudy (mathstudent55):

Now we need to pick from choices 2 and 4, which one also has the correct second equation.

OpenStudy (mathstudent55):

The second equation comes from this statement: "Variable c is also 3 less than variable w. " The equation for this statement is: c = w - 3 Can you pick the correct choice from choices 2 and 4?

OpenStudy (anonymous):

2

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

thank you so much :)